# Rare and exclusive few-body decays of the Higgs, Z, W bosons, and the top quark

### v3

@@ II.3 Input parameters of the calculations
Table 2: Numerical values of SM parameters used here in theoretical calculations of various branching fractions. Most of the values are from the PDG [58] (the leptonium masses, $m_{\ell\ell}$, are given as twice the single lepton masses, ignoring tiny binding energies), except for those where a specific reference is provided.
@@
$m_{\mathrm{ee}}$ $1.02991$ MeV $\alpha(0)$ 1/137.036 $m_{\mathrm{\pi^{\pm}}}$ 0.13957 $f_{\mathrm{\pi}}$ 0.1304 [44]
$m_{\mu\mu}$ 0.21126 GeV $\alpha(m_{\mathrm{Z}})$ 1/128.943 $m_{\mathrm{K^{\pm}}}$ 0.493677 $f_{\mathrm{K}}$ 0.1562 [44]
$m_{\tau\tau}$ 3.5537 GeV $\alpha_{\text{s}}(m_{{}_{\mathrm{Z}}})$ 0.1180 $\lambda_{\mathrm{K}}$ 0.255 [67]
$m_{\mathrm{c}}$ 1.50 GeV $\sin^{2}{\theta_{\mathrm{w}}}$ 0.2351 $m_{\mathrm{\rho^{\pm}}}$ 0.77526 $f_{\mathrm{\rho}}$ 0.212 [44]
$m_{\mathrm{b}}$ 4.18 GeV $\rm G_{F}$ $1.1664\cdot 10^{-5}$ GeV-2 $m_{\mathrm{K^{*\pm}}}$ 0.89166 $f_{\mathrm{K^{*}}}$ 0.203 [44]
$m_{\mathrm{t}}$ 172.69 GeV $|V_{\mathrm{ud,cs}}|$ 0.974 $m_{\mathrm{\phi}}$ 1.0194 $f_{\mathrm{\phi}}$ 0.223 [14]
$m_{\mathrm{W}}$ 80.377 GeV $|V_{\mathrm{us,cd}}|$ 0.225 $m_{\mathrm{D^{\pm}}}$ 1.86966 $f_{\mathrm{D}}$ 0.212 [68]
$m_{\mathrm{Z}}$ 91.188 GeV $|V_{\mathrm{cb,ts}}|$ 0.041 $\lambda_{\mathrm{D}}$ 0.354 [69]
$m_{\mathrm{H}}$ 125.25 GeV $|V_{\mathrm{ub}}|$ 0.00382 $m_{\mathrm{D_{s}^{\pm}}}$ 1.96835 $f_{\mathrm{D_{s}}}$ 0.2499 [68]
$\Gamma_{\mathrm{W}}$ 2.085 GeV $|V_{\mathrm{tb}}|$ 0.999 $m_{\mathrm{D^{*0}}}$ 2.007 $f_{\mathrm{D^{*0}}}$ 1.28 $f_{\mathrm{D}}$ [70]
$\Gamma_{\mathrm{W\to\mu\nu}}$ $0.1063~{}\Gamma_{\mathrm{W}}$ $|V_{\mathrm{td}}|$ 0.0086 $m_{\mathrm{D^{*\pm}}}$ 2.01027 $f_{\mathrm{D^{*}}}$ 1.28 $f_{\mathrm{D}}$ [70]
$\Gamma_{\mathrm{Z}}$ 2.4955 GeV $m_{\mathrm{D_{s}^{*\pm}}}$ 2.1123 $f_{\mathrm{D_{s}^{*}}}$ 1.26 $f_{\mathrm{D_{s}}}$ [70]
$\Gamma_{\mathrm{Z\to\ell\ell}}$ $0.0337~{}\Gamma_{\mathrm{Z}}$ $m_{\mathrm{B^{\pm}}}$ 5.27934 $f_{\mathrm{B}}$ 0.190 [68]
$\Gamma_{\mathrm{H}}$ 4.1 MeV [71] $\lambda_{\mathrm{B}}$ 0.338 [72]
$\Gamma_{\mathrm{H\to\gamma\gamma}}$ $2.5\cdot 10^{-3}~{}\Gamma_{\mathrm{H}}$ $m_{\mathrm{B^{*\pm,0}}}$ 5.32471 $f_{\mathrm{B^{*}}}$ 0.941 $f_{\mathrm{B}}$ [73]
$\Gamma_{\mathrm{t}}$ 1.331 GeV [74] $f_{\mathrm{B_{\mathrm{s}}}}$ 0.2303 [68]
$\lambda_{\mathrm{B_{s}}}$ 0.438 [75]
$m_{\mathrm{B_{s}^{*0}}}$ 5.415 $f_{\mathrm{B_{\mathrm{s}}^{*}}}$ 0.953 $f_{\mathrm{B_{\mathrm{s}}}}$ [73]
$m_{\mathrm{B_{\mathrm{c}}^{\pm}}}$ 6.27447 $f_{\mathrm{B_{\mathrm{c}}}}$ 0.427 [73]
$m_{\mathrm{B_{\mathrm{c}}^{*\pm}}}$ 6.32877 [73] $f_{\mathrm{B_{\mathrm{c}}^{*}}}$ 0.988 $f_{\mathrm{B_{\mathrm{c}}}}$ [73]
@@ III.1 Rare two-, three-, and four-body Higgs boson decays
Figure 2 shows rare decays of the H boson into two neutrinos, a photon plus two neutrinos, three or four photons, and a Z boson plus two gluons or two photons. Calculations of rates for these processes, which proceed via virtual loops in the SM, do not exist in many cases to our knowledge. Since they are very suppressed and have no obvious phenomenological impact, they are not included in the standard Higgs decay codes [71], although they are intriguing for diverse reasons exposed below. We have therefore computed them with the MadGraph5_aMC@NLO program [76] (called MG5_AMC, hereafter) with QCD and EW loop corrections up to NLO accuracy [77, 78], including massive fermions in the loops (where needed), using the input parameters of Table 2. The corresponding results are listed in Table 3.
@@
Our first calculation is that of the invisible two-body Higgs boson decay into a neutrino-antineutrino pair, which is infinitesimal in the SM ($\mathcal{B}\approx 10^{-36}$) compared to the standard invisible (four-neutrino) $\rm H\to ZZ^{\star}\to 4\nu$ decay ($\mathcal{B}\approx 0.1\%$[71]. However, since the SM assumption of massless $\nu$’s is incorrect, the $\rm H\to\nu\overline{\nu}$ decay can receive extra contributions depending on the mechanism of neutrino mass generation actually realized in nature. Therefore, it is a process worth keeping an eye on, when considering invisible Higgs decays and accounting for massive neutrinos. The second diagram of Fig. 2 shows the photon-plus-neutrinos decay, which proceeds through Z and W loops and, since the neutrino pair goes undetected, it appears experimentally as an unbalanced monophoton decay of the Higgs boson. Such a final state is shared by many exotic BSM Higgs decays [9], and it is worth computing its SM rate. This decay is actually the least rare considered in this whole survey, and has a branching fraction of $\mathcal{B}=3.74\cdot 10^{-4}$ that is about 20% larger than the naive estimate given by $\mathcal{B}(\rm H\to Z\gamma)=1.54\cdot 10^{-3}$ [71] combined with $\mathcal{B}(\rm Z\to\nu\overline{\nu})=0.20$ [58]. This is so because extra W-induced channels (not shown in Fig. 2) contribute to the amplitude.
The third ultrarare Higgs decay of interest here is $\rm H\to 3\gamma$. At variance with the naive assumption that it is forbidden for a scalar particle to decay into three photons, because such a process would violate charge-conjugation (C) symmetry, $\rm C(H)=0\neq C(3\gamma)=(-1)^{3}=-1$, such a decay is possible, albeit extremely suppressed, for EW-induced decays. In a situation akin to the case of the triphoton decay of the scalar positronium state (para-positronium) [79, 80] or of the $\pi^{0}$ meson [81], which is forbidden in QED but not in the EW theory, such a Higgs decay can proceed through the W box shown in the top-right panel of Fig. 2. Since $\rm H\to 3\gamma$ violates C-symmetry, it must also violate parity in order to conserve CP and, therefore, the final state must be composed of three spatially symmetric photons with vanishing total angular momentum. As a consequence, the partial width of this channel features an utterly small $\mathcal{O}(10^{-40})$ probability. The box diagram is however not the only contribution to this decay, which can also proceed via $\rm H\to Z(\gamma\gamma)\gamma$. Estimates of this latter contribution are of the same size as the box ones [82].
Table 3: Compilation of rare Higgs decays to two neutrinos, a photon plus two neutrinos, three or four photons, and a Z boson plus two photons or two gluons. For each decay, we provide the theoretical branching fraction computed with MG5_aMC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-hh.
@@
Branching fraction Framework 2024 HL-LHC FCC-ee FCC-hh
$\rm H\,\to$ $\rm\nu+\overline{\nu}$  $7.2$$\times~{}10^{-36}$ NLO SM, MG5_aMC (this work)
$\rm\gamma+\nu+\overline{\nu}$  $3.4$$\times~{}10^{-4}$ NLO SM [83]
 $3.6$$\times~{}10^{-4}$ NLO SM, MG5_aMC (this work)
$\rm\gamma+\gamma+\gamma~{}(box)$  $1.0$$\times~{}10^{-40}$ NLO SM, MG5_aMC (this work)
$\rm\gamma+\gamma+\gamma+\gamma$  $5.4$$\times~{}10^{-12}$ NLO SM, MG5_aMC (this work)
$\rm Z+\gamma+\gamma$  $2.0$$\times~{}10^{-9}$ NLO SM(with cuts) [84]
 $2.4$$\times~{}10^{-9}$ NLO SM, MG5_aMC (this work)
$\rm Z+g+g$  $3.6$$\times~{}10^{-7}$ NLO SM [85]
 $5.5$$\times~{}10^{-7}$ NLO SM [^2] With updated SM parameters [86]
 $6.3$$\times~{}10^{-7}$ NLO SM, MG5_aMC (this work)
@@
Two other Higgs boson decays into three gauge bosons are possible: $\rm H\to Zgg$ and $\rm H\to Z\gamma\gamma$. They have been considered before in the literature, albeit in the $m_{\mathrm{t}}\to\infty$ limit and/or with old SM parameters [86, 84, 85]. Both decays proceed through a fermion box (mostly, a top quark box) as shown in the bottom of Fig. 2 (left and center diagrams). The $\rm H\to Z\gamma\gamma$ decay rate is much more suppressed than the naive expectation of it being similar to $\mathcal{B}(\rm H\to Z\gamma)\approx 1.5\cdot 10^{-3}$ times the QED coupling for the extra photon emission because, due to the charge conjugation properties of the gauge boson couplings, there are no W bosons in the virtual box. It is worth noting that the similar box diagram of the $\rm H\to\gamma gg$ decay, which results by changing the Z boson by an emitted photon, is forbidden by the generalized Furry’s theorem [87] that applies for the photon case but is evaded by the axial coupling of the Z boson to fermions in the $\rm H\to Zgg$ case (this decay can however go through $\rm H\to Z(gg)\gamma$ with an estimated rate of $\mathcal{O}(10^{-40})$ [82]). We have calculated both processes here with MG5_aMC –including top-, bottom-quarks, and tau leptons in the box– finding branching fractions of $\mathcal{B}(\rm H\to Zgg)=6.3\cdot 10^{-7}$ and $\mathcal{B}(\rm H\to Z\gamma\gamma)=2.4\cdot 10^{-9}$, respectively. With such low $\mathcal{B}$ values, one expects short of one $\rm H\to Zgg$ event at the FCC-ee, and both very rare channels are only producible at a machine such as FCC-hh, where they offer a relatively clean signal of two energetic gluons or photons plus a back-to-back lepton-antilepton pair.
@@
Figure 3: Theoretical decay branching fractions of the Higgs boson into rare two-, three-, or four- gauge bosons and/or neutrinos, computed in this work shown as a function of the Higgs boson mass (left) and as blue bars in negative log scale (right). In the left panel, the dashed lines for $m_{\mathrm{H}}<m_{\mathrm{Z}}$ show the $\rm H\to Z*gg,\,Z*\gamma\gamma$ decay rates with offshell Z bosons. In the right panel the red vertical lines indicate the FCC-ee (solid) and FCC-hh (dashed) reach based only on the total number of H bosons to be produced at both facilities.
@@
Finally, it is interesting to compute the 4-photon decay of the Higgs boson as it may constitute a background for exotic decays into a pair of light axion-like or scalar particles, each of which further decays into two photons, $\mathrm{H}\to a(\gamma\gamma)\,a(\gamma\gamma)$. The Higgs $4\gamma$ decay has a branching fraction of $\mathcal{B}=4.56\cdot 10^{-12}$, which is 28 orders-of-magnitude larger than the 3-$\gamma$ box-mediated decay, as C and P conservation is not an issue here, and the rate is only suppressed due to the presence of heavy charged-particle loops. Among all rare channels discussed in this section, LHC searches have been performed to date for the $4\gamma$ final state alone [66, 88, 89, 90] (searches for $3\gamma$ decays have focused on Z’ resonances off the Higgs peak [66]), but unfortunately, the limits have been set only on the process $\rm H\to a(\gamma\gamma)a(\gamma\gamma)$ with two intermediate ALPs decaying into photons. It would not be too difficult for ATLAS/CMS to recast these searches into upper bounds on the $\rm H\to 4\gamma$ “continuum” decay. For reference, Fig. 3 shows the branching fractions for the six very rare decays of the Higgs boson into two-, three-, and four particles, computed here, as a function of the Higgs boson mass $m_{\mathrm{H}}$ (left), and in negative log scale for the SM Higgs mass (right).

### v4

@@ II.3 Input parameters of the calculations
Table 2: Numerical values of SM parameters used here in theoretical calculations of various branching fractions. Most of the values are from the PDG [ParticleDataGroup:2024cfk] (the leptonium masses, $m_{\ell\ell}$, are given as twice the single lepton masses, ignoring tiny binding energies), except for those where a specific reference is provided.
@@
$m_{\mathrm{ee}}$ $1.02991$ MeV $\alpha(0)$ 1/137.036 $m_{\mathrm{\pi^{\pm}}}$ 0.13957 $f_{\mathrm{\pi}}$ 0.1304 [Grossman:2015cak]
$m_{\mu\mu}$ 0.21126 GeV $\alpha(m_{\mathrm{Z}})$ 1/128.943 $m_{\mathrm{K^{\pm}}}$ 0.493677 $f_{\mathrm{K}}$ 0.1562 [Grossman:2015cak]
$m_{\tau\tau}$ 3.5537 GeV $\alpha_{\text{s}}(m_{{}_{\mathrm{Z}}})$ 0.1180 $\lambda_{\mathrm{K}}$ 0.255 [Puhan:2023ekt]
$m_{\mathrm{c}}$ 1.50 GeV $\sin^{2}{\theta_{\mathrm{w}}}$ 0.2351 $m_{\mathrm{\rho^{\pm}}}$ 0.77526 $f_{\mathrm{\rho}}$ 0.212 [Grossman:2015cak]
$m_{\mathrm{b}}$ 4.18 GeV $\rm G_{F}$ $1.1664\cdot 10^{-5}$ GeV-2 $m_{\mathrm{K^{*\pm}}}$ 0.89166 $f_{\mathrm{K^{*}}}$ 0.203 [Grossman:2015cak]
$m_{\mathrm{t}}$ 172.69 GeV $|V_{\mathrm{ud,cs}}|$ 0.974 $m_{\mathrm{\phi}}$ 1.0194 $f_{\mathrm{\phi}}$ 0.223 [Konig:2015qat]
$m_{\mathrm{W}}$ 80.377 GeV $|V_{\mathrm{us,cd}}|$ 0.225 $m_{\mathrm{D^{\pm}}}$ 1.86966 $f_{\mathrm{D}}$ 0.212 [FlavourLatticeAveragingGroupFLAG:2021npn]
$m_{\mathrm{Z}}$ 91.188 GeV $|V_{\mathrm{cb,ts}}|$ 0.041 $\lambda_{\mathrm{D}}$ 0.354 [Lu:2021ttf]
$m_{\mathrm{H}}$ 125.25 GeV $|V_{\mathrm{ub}}|$ 0.00382 $m_{\mathrm{D_{s}^{\pm}}}$ 1.96835 $f_{\mathrm{D_{s}}}$ 0.2499 [FlavourLatticeAveragingGroupFLAG:2021npn]
$\Gamma_{\mathrm{W}}$ 2.085 GeV $|V_{\mathrm{tb}}|$ 0.999 $m_{\mathrm{D^{*0}}}$ 2.007 $f_{\mathrm{D^{*0}}}$ 1.28 $f_{\mathrm{D}}$ [Becirevic:2012ti]
$\Gamma_{\mathrm{W\to\mu\nu}}$ $0.1063~\Gamma_{\mathrm{W}}$ $|V_{\mathrm{td}}|$ 0.0086 $m_{\mathrm{D^{*\pm}}}$ 2.01027 $f_{\mathrm{D^{*}}}$ 1.28 $f_{\mathrm{D}}$ [Becirevic:2012ti]
$\Gamma_{\mathrm{Z}}$ 2.4955 GeV $m_{\mathrm{D_{s}^{*\pm}}}$ 2.1123 $f_{\mathrm{D_{s}^{*}}}$ 1.26 $f_{\mathrm{D_{s}}}$ [Becirevic:2012ti]
$\Gamma_{\mathrm{Z\to\ell\ell}}$ $0.0337~\Gamma_{\mathrm{Z}}$ $m_{\mathrm{B^{\pm}}}$ 5.27934 $f_{\mathrm{B}}$ 0.190 [FlavourLatticeAveragingGroupFLAG:2021npn]
$\Gamma_{\mathrm{H}}$ 4.1 MeV [Djouadi:2018xqq] $\lambda_{\mathrm{B}}$ 0.338 [Mandal:2023lhp]
$\Gamma_{\mathrm{H\to\gamma\gamma}}$ $2.5\cdot 10^{-3}~\Gamma_{\mathrm{H}}$ $m_{\mathrm{B^{*\pm,0}}}$ 5.32471 $f_{\mathrm{B^{*}}}$ 0.941 $f_{\mathrm{B}}$ [Colquhoun:2015oha]
$\Gamma_{\mathrm{t}}$ 1.331 GeV [Chen:2022wit] $f_{\mathrm{B_{\mathrm{s}}}}$ 0.2303 [FlavourLatticeAveragingGroupFLAG:2021npn]
$\lambda_{\mathrm{B_{s}}}$ 0.438 [Khodjamirian:2020hob]
$m_{\mathrm{B_{s}^{*0}}}$ 5.415 $f_{\mathrm{B_{\mathrm{s}}^{*}}}$ 0.953 $f_{\mathrm{B_{\mathrm{s}}}}$ [Colquhoun:2015oha]
$m_{\mathrm{B_{\mathrm{c}}^{\pm}}}$ 6.27447 $f_{\mathrm{B_{\mathrm{c}}}}$ 0.427 [Colquhoun:2015oha]
$m_{\mathrm{B_{\mathrm{c}}^{*\pm}}}$ 6.32877 [Colquhoun:2015oha] $f_{\mathrm{B_{\mathrm{c}}^{*}}}$ 0.988 $f_{\mathrm{B_{\mathrm{c}}}}$ [Colquhoun:2015oha]
@@ III.1 Rare two-, three-, and four-body Higgs boson decays
Figure 2 shows rare decays of the H boson into two neutrinos, a photon plus two neutrinos, three or four photons, and a Z boson plus two gluons or two photons. Calculations of rates for these processes, which proceed via virtual loops in the SM, do not exist in many cases to our knowledge. Since they are very suppressed and have no obvious phenomenological impact, they are not included in the standard Higgs decay codes [Djouadi:2018xqq], although they are intriguing for diverse reasons exposed below. We have therefore computed them with the MadGraph5_aMC@NLO program [Alwall:2014hca] (called MG5_AMC, hereafter) with QCD and EW loop corrections up to NLO accuracy [Hirschi:2015iia, Frederix:2018nkq], including massive fermions in the loops (where needed), using the input parameters of Table 2. The corresponding results are listed in Table 3.
@@
Our first calculation is that of the invisible two-body Higgs boson decay into a neutrino-antineutrino pair, which is infinitesimal in the SM ($\mathcal{B}\approx 10^{-26}$) compared to the standard invisible (four-neutrino) $\rm H\to ZZ^{\star}\to 4\nu$ decay ($\mathcal{B}\approx 0.1\%$[Djouadi:2018xqq]. However, since the SM assumption of massless $\nu$’s is incorrect, the $\rm H\to\nu\overline{\nu}$ decay can receive extra contributions depending on the mechanism of neutrino mass generation actually realized in nature. Therefore, it is a process worth keeping an eye on, when considering invisible Higgs decays and accounting for massive neutrinos [dEnterria:2025rjj]. The second diagram of Fig. 2 shows the photon-plus-neutrinos decay, which proceeds through Z and W loops and, since the neutrino pair goes undetected, it appears experimentally as an unbalanced monophoton decay of the Higgs boson. Such a final state is shared by many exotic BSM Higgs decays [Curtin:2013fra], and it is worth computing its SM rate. This decay is actually the least rare considered in this whole survey, and has a branching fraction of $\mathcal{B}=3.74\cdot 10^{-4}$ that is about 20% larger than the naive estimate given by $\mathcal{B}(\rm H\to Z\gamma)=1.54\cdot 10^{-3}$ [Djouadi:2018xqq] combined with $\mathcal{B}(\rm Z\to\nu\overline{\nu})=0.20$ [ParticleDataGroup:2024cfk]. This is so because extra W-induced channels (not shown in Fig. 2) contribute to the amplitude.
The third ultrarare Higgs decay of interest here is $\rm H\to 3\gamma$. At variance with the naive assumption that it is forbidden for a scalar particle to decay into three photons, because such a process would violate charge-conjugation (C) symmetry, $\rm C(H)=+1\neq C(3\gamma)=(-1)^{3}=-1$, such a decay is possible, albeit extremely suppressed, for EW-induced decays. In a situation akin to the case of the triphoton decay of the scalar positronium state (para-positronium) [Bernreuther:1981ah, Pokraka:2017ore] or of the $\pi^{0}$ meson [Dicus:1975cz], which is forbidden in QED but not in the EW theory, such a Higgs decay can proceed through the W box shown in the top-right panel of Fig. 2. Since $\rm H\to 3\gamma$ violates C-symmetry, it must also violate parity in order to conserve CP and, therefore, the final state must be composed of three spatially symmetric photons with vanishing total angular momentum. As a consequence, the partial width of this channel features an utterly small $\mathcal{O}(10^{-40})$ probability. The box diagram is however not the only contribution to this decay, which can also proceed via $\rm H\to Z(\gamma\gamma)\gamma$. Estimates of this latter contribution are of the same size as the box ones [dEnterria:2025rjj].
Table 3: Compilation of rare Higgs decays to two neutrinos, a photon plus two neutrinos, three or four photons, and a Z boson plus two photons or two gluons. For each decay, we provide the theoretical branching fraction computed with MG5_aMC. The last two columns indicate whether the decay can be produced at FCC-ee/FCC-$h$.
@@
Branching fraction Framework 2024 HL-LHC FCC-$e$ FCC-$h$
$\rm H\,\to$ $\rm\nu+\overline{\nu}$ $2.0$ $\times~10^{-26}$ loop SM, MG5_aMC (this work)
$\rm\gamma+\nu+\overline{\nu}$ $3.4$ $\times~10^{-4}$ loop SM [Sun:2013cba]
$3.6$ $\times~10^{-4}$ loop SM, MG5_aMC (this work)
$\rm\gamma+\gamma+\gamma$~ $1.0$ $\times~10^{-40}$ loop SM, MG5_aMC (this work)
$\rm\gamma+\gamma+\gamma+\gamma$ $5.4$ $\times~10^{-12}$ loop SM, MG5_aMC (this work)
$\rm Z+\gamma+\gamma$ $2.0$ $\times~10^{-9}$ loop SM(with cuts) [Abbasabadi:2004wq]
$2.4$ $\times~10^{-9}$ loop SM, MG5_aMC (this work)
$\rm Z+g+g$ $3.6$ $\times~10^{-7}$ loop SM [Abbasabadi:2008zz]
$5.5$ $\times~10^{-7}$ loop SM [^2] With updated SM parameters [Kniehl:1990yb]
$6.3$ $\times~10^{-7}$ loop SM, MG5_aMC (this work)
@@
Two other Higgs boson decays into three gauge bosons are possible: $\rm H\to Zgg$ and $\rm H\to Z\gamma\gamma$. They have been considered before in the literature, albeit in the $m_{\mathrm{t}}\to\infty$ limit and/or with old SM parameters [Kniehl:1990yb, Abbasabadi:2004wq, Abbasabadi:2008zz]. Both decays proceed through a fermion box (mostly, a top quark box) as shown in the bottom of Fig. 2 (left and center diagrams). The $\rm H\to Z\gamma\gamma$ decay rate is much more suppressed than the naive expectation of it being similar to $\mathcal{B}(\rm H\to Z\gamma)\approx 1.5\cdot 10^{-3}$ times the QED coupling for the extra photon emission because, due to the charge conjugation properties of the gauge boson couplings, there are no W bosons in the virtual box. It is worth noting that the similar box diagram of the $\rm H\to\gamma gg$ decay, which results by changing the Z boson by an emitted photon, is forbidden by the generalized Furry’s theorem [Nishijima:1951] that applies for the photon case but is evaded by the axial coupling of the Z boson to fermions in the $\rm H\to Zgg$ case (this decay can however go through $\rm H\to Z(gg)\gamma$ with an estimated rate of $\mathcal{O}(10^{-40})$ [dEnterria:2025rjj]). We have calculated both processes here with MG5_aMC –including top-, bottom-quarks, and tau leptons in the box– finding branching fractions of $\mathcal{B}(\rm H\to Zgg)=6.3\cdot 10^{-7}$ and $\mathcal{B}(\rm H\to Z\gamma\gamma)=2.4\cdot 10^{-9}$, respectively. With such low $\mathcal{B}$ values, one expects short of one $\rm H\to Zgg$ event at the FCC-ee, and both very rare channels are only producible at a machine such as FCC-hh, where they offer a relatively clean signal of two energetic gluons or photons plus a back-to-back lepton-antilepton pair.
@@
Figure 3: Theoretical decay branching fractions of the Higgs boson into rare two-, three-, or four- gauge bosons and/or neutrinos, computed in this work shown as a function of the Higgs boson mass (left) and as blue bars in negative log scale (right). In the left panel, the dashed lines for $m_{\mathrm{H}}<m_{\mathrm{Z}}$ show the $\rm H\to Z^{*}gg,\,Z^{*}\gamma\gamma$ decay rates with offshell Z bosons. In the right panel the red vertical lines indicate the FCC-ee (solid) and FCC-hh (dashed) reach based only on the total number of H bosons to be produced at both facilities.
@@
Finally, it is interesting to compute the 4-photon decay of the Higgs boson as it may constitute a background for exotic decays into a pair of light axion-like or scalar particles, each of which further decays into two photons, $\mathrm{H}\to a(\gamma\gamma)\,a(\gamma\gamma)$. The Higgs $4\gamma$ decay has a branching fraction of $\mathcal{B}=4.56\cdot 10^{-12}$, which is 28 orders-of-magnitude larger than the 3-$\gamma$ box-mediated decay, as C and P conservation is not an issue here, and the rate is only suppressed due to the presence of heavy charged-particle loops. Among all rare channels discussed in this section, LHC searches have been performed to date for the $4\gamma$ final state alone [ATLAS:2015rsn, CMS:2022fyt, CMS:2022xxa, ATLAS:2023ian] (searches for $3\gamma$ decays have focused on Z’ resonances off the Higgs peak [ATLAS:2015rsn]), but unfortunately, the limits have been set only on the process $\rm H\to a(\gamma\gamma)a(\gamma\gamma)$ with two intermediate ALPs decaying into photons. It would not be too difficult for ATLAS/CMS to recast these searches into upper bounds on the $\rm H\to 4\gamma$ “continuum” decay. For reference, Fig. 3 shows the branching fractions for the six very rare decays of the Higgs boson into two-, three-, and four particles, computed here, as a function of the Higgs boson mass $m_{\mathrm{H}}$ (left), and of negative log branching fraction for the SM Higgs mass (right).